OpenStudy (anonymous):

Find the area of a parallelogram if a base and corresponding altitude have the indicated lengths. Base 1 1/2 feet, altitude 6 inches.

OpenStudy (anonymous):

area = 3/2 * 1/2 = 3/4 sq ft

OpenStudy (anonymous):

ok cherri?